Dave can dig a ditch in 6 hours, and Sam can dig a ditch in 4 hours. Frank can fill a ditch in 8 hours. If Dave and Sam dig a ditch while Frank tries to fill it, how long does it take to dig the ditch altogether?
Answer by mananth(16949)

Dave can dig aditch in6 hours
So he can do 1/6 of job in 1 hour
.
Sam can dig in 4 hours.
so he can do 1/4 of the job in 1 hour
.
Frank fills the ditch in 8 hours
So he can do 1/8 of the job of filling in 1 hour
.
1/6 + 1/4 of the job of digging they do in 1 hour
=2+3 /12
= 5/12
.
5/12 - 1/8 is the effective digging in 1 hour
10-3 / 24
=7/24 is the effective digging in 1 hour
So to complete the job of digging they will take 24/7 hours
= 3.42 hours